NEW HIGH RESOLUTION ANALYSIS OF THE ν_{8}+ν_{9} BAND OF H^{14}NO_{3} BY FOURIER TRANSFORM SPECTROSCOPY

\underline{A. PERRIN}, J.-M. FLAUD, Laboratoire de Photophysique Moléculaire, CNRS, Université Paris Sud, Campus d’Orsay, Bat 210, 91405 Orsay Cedex, France; A. GOLDMAN, R.D. BLATHERWICK, and F.J. MURCRAY, University of Denver, Physics Department, 2112 E.Wesley Ave. Denver, CO 80208 USA; C.P. RINSLAND, Atmospheric Science Division, NASA Langley Research Center, Hampton, VA 23681-001, USA.

Using new Fourier transform spectra recorded at high resolution at Denver, in the 1100-1300cm^{−1} spectral range, it has been possible to perform a new and more extended study of the ν_{8}+ν_{9} band of HNO_{3}. ^{a} ^{b} Actually, more than 4400 lines belonging to the ν_{8}+ν_{9} band could be assigned up to high J and K_{a} values (J≤75 and K_{a}≤45), leading to the determination of 1700 infrared energy levels of the v_{8}=v_{9}=1 state. As in the recent infrared analysis of the ν_{8}+ν_{9} band performed by Wang^{b} the Fermi-type resonance coupling the rotational levels of the v_{8}=v_{9}=1 vibrational state with those of the v_{6}=v_{7}=1 ”dark state” was explicitly taken into account during the energy levels calculations. This allowed to reproduce the infrared levels to within their experimental accuracy. However, it should be pointed out that when dealing with the microwave measurements ^{c} within ν_{8}+ν_{9} , some microwave lines could not be calculated properly. Finally the determination of the A- and B-type components of the transition moment operator of the ν_{8}+ν_{9} band was performed from a least squares fit calculation performed on experimental line intensities measured in the spectrum.

aA. Maki, J. Mol. Spectrosc. \underline{136}, 105 (1989).

bW. F. Wang, P. P. Ong, E.C. Looi and H.H. Teo, J. Mol. Spectrosc. \underline{183}, 407 (1997).

cT. M. Goyette and F.C. De Lucia, J. Mol. Spectrosc. \underline{139}, 241 (1990).